OpenStudy (anonymous):

The equation below shows the perimeter f(x) of several rectangular pieces of cardboard having width x inches: f(x) = 52 + 2x The perimeter of each piece is less than 100 inches. Which of the following best describes the domain of f(x)? a) all real numbers greater than 0 and less than 100 b)all real numbers greater than 0 and less than 24 c)all integers greater than 0 and less than 100 d)all integers greater than 0 and less than 24

OpenStudy (anonymous):

Your idea?

OpenStudy (anonymous):

i dont have an idea ..

OpenStudy (anonymous):

What is the perimeter formula of a rectangular in general?

OpenStudy (anonymous):

l x h x w

OpenStudy (anonymous):

nope, 2(l+w)

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hence, your f(x) = 52 + 2x = 2( 26 + x) ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, just let it < 100 , solve for x, dat sit

OpenStudy (anonymous):

just solve one to the right , I mean 52 + 2x < 100 for x and x is the width of rectangular, it must be >0, and 26 is the length of the rectangular and it is an integer, that forces x must be an integer also.

OpenStudy (anonymous):

so would it be c or d?

OpenStudy (anonymous):

c